OpenStudy (anonymous):

Use the first derivative to find all critical points and use the second derivative to find all the inflection points. f(x)=2x^3+3x^2-36x+9 The critical point at x= _____ The critical point at x= _____ The inflection point is x= _____

OpenStudy (anonymous):

I found: f'(x) = 6x^2+6x-36 f"(x) = 12x+6 But I'm not sure what to do with f' since it can only be factored to 6(x^2+x-6) The inflection point is x=-2, I think? 12x+6 = 0 12x = -6 x = -2

OpenStudy (anonymous):

Oh yeah, duh. I would have figured that out once I plugged it in. But I still need help with the critical points.

OpenStudy (ranga):

f'(x) = 6(x^2+x-6) f'(x) = 6(x+3)(x-2) To find critical points, set f'(x) = 0 and solve for x 6(x+3)(x-2) = 0 x = -3, +2 are critical points.

OpenStudy (ranga):

Point of inflection at x = -1/2.

OpenStudy (anonymous):

Thanks!

OpenStudy (ranga):

you are welcome.